OpenStudy (anonymous):

Determine whether y varies directly with x. If so, find the constant of variation k and write the equation.

OpenStudy (anonymous):

Is there some sort of table, graph, etc. that you are not giving us?

OpenStudy (anonymous):

The specific problem is 2x-3y=6.

OpenStudy (anonymous):

In order for y to vary directly with x, y and x must always have the same ratio (y / x = k): When x = 3: 2(3) - 3y = 6 6 - 3y = 6 6 = 3y + 6 0 = 3y y = 0 k = 0 / 3 = 0 When x = 6: 2(6) - 3y = 6 12 - 3y = 6 12 = 6 + 3y 6 = 3y y = 2 k = 2 / 3 As we can see, y and x do not always have the same ratio, so y does not vary directly with x.

OpenStudy (anonymous):

Hope that helps!

OpenStudy (anonymous):

Thank you!! I was trying to figure it out for a while

OpenStudy (anonymous):

A medal would be nice :).